What should be the radius of a space station spinning with an angular speed of 2 rpm such that an astronaut inside the station experiences artificial gravity with g = 10 ms-2?
1
Expert's answer
2017-03-03T11:49:05-0500
F=m*g; gravity force
F=m*w^2*R; centrifugal force

m*g=m*w^2*R

w=2 rpm
g=10 m/s

R=(m*g)/(m*w^2)=g/w^2=10/(2/60)^2=10*300=9000 m

Answer:
radius of a space station should be - 9000 m